OpenStudy (anonymous):

Solve. w minus four point eight equals five point one w equals nine point nine w equals five point five eight w equals 3 w equals zero point three

OpenStudy (md152727):

w-4.8=5.1 5.1-4.8= 0.3 so D

OpenStudy (agreene):

start by writing it out in numbers: \[w - 4.8 = 5.1\]